Crew 190 Sol Summary Report
Phoenix awaken! But still a bit buzzy.
Michael Saint-Guillain (XO)
Ready to continue simulation on Sol 6, 08:00
Sol Activity Summary:
8:15 Medic inspection: memory and reflex tests
9:30 EVA: Bastien (cartography – 3D mapping), Martin (experiment check), Sophie (muography), Fred (photos)
9:30 Others: Scientific work
11:00 EVA ended; Crew meeting before lunch
14:30 Scientific work: muography (Sophie), video extraction & compilation (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), Phoenix troubleshooting (Michael)
18:00 Report redaction
Notes: EVA (Max, Mario, Ariane, Mick) in the morning (scientific work for the others). Scientific work in the afternoon.
Anomalies in work:
Today we tried to get the Phoenix remote controlled rover working; after fixing some battery issues, we noticed that the rover was not able to connect to the (dedicated) router wifi. We had to manually change the clock of the rover it self in order to fix that problem: the rover had the correct date, but not the router, which we cannot fix (probably the internal clock battery of the router is dead). With similar date/times, router and rover are now able to talk to each other. The video transmission is working, but the commands are still bugging (can’t move the wheels, and cams movement is working randomly).
Sunny in the morning. Cloudy in the afternoon. Nice weather.
Crew Physical Status:
See EVA report.
Reports to be file:
EVA#7 request for Sol6